Anna is thinking about making an improvement to her home. This modification will cost $10,000, but it will also increase the value of the home by $10,000 when Anna plans to sell it in 10 years. The market interest rate is 10%. 


What is the minimum amount at which Anna must personally value the improvement to be willing to go through with it? 


1
Expert's answer
2021-11-14T17:23:16-0500

Solution:

Derive the future value:

FV = PMT "\times" (1 + r)n

PMT = 10,000

r = 10%

n = 10 years


FV = 10,000 "\times" (1 + 0.1)10 = 25,937

FV = $25,937


Anna must personally value the improvement at $25,937 to be willing to go through with it